# C-LoRA: Contextual Low-Rank Adaptation for Uncertainty Estimation in Large Language Models

### v2

@@ 5 Related Works
LLMs have demonstrated remarkable successes across a wide range of applications, including code generation, scientific reasoning, and open-domain question answering; however, they are also notorious for hallucination—a phenomenon in which the model generates fluent but factually incorrect or unsupported content that is not grounded in the input or external knowledge. To mitigate hallucination, a variety of strategies have been proposed, including Reinforcement Learning from Human Feedback (RLHF) to align model outputs with human preferences (liang2024learning, ; wen2024policy, ), and Retrieval-Augmented Generation (RAG) to ground responses in external knowledge (lewis2020retrieval, ; gao2023retrieval, ). Additionally, probabilistic methods such as BNNs (wang2024blobbayesianlowrankadaptation, ; yang2024bayesianlowrankadaptationlarge, ) and conformal prediction (kumar2023conformal, ; quach2023conformal, ) have been demonstrated to be able to quantify uncertainty and flag unreliable outputs. Recent work also shows the effectiveness of prompt-level interventions (jiang2023calibrating, ; yadkori2024believe, ) by carefully crafting instructions and uncertainty-aware prompts to reduce hallucinated content without altering the model architecture.
@@ 7 Limitations
Evaluating uncertainty in language models remains an open challenge due to the lack of standardized benchmarks with ground-truth uncertainty labels. While our evaluation—based on held-out test sets and established metrics like ECE and NLL—provides meaningful insight into model calibration and predictive confidence, these metrics offer only an indirect view of uncertainty quality. Developing task-specific benchmarks or human-in-the-loop protocols for evaluating uncertainty in NLP remains an important direction for future work.
@@ Appendix E On the Importance of Flexibility
Table E summarizes the performance of DE and FE with M=0 and M=10 under in-distribution scenario for the six common-sense reasoning tasks. These models are ordered from top to bottom for each metric, such that each successive model is more flexible than the previous one. From Table E, it is clear that as the flexibility increases, the performance generally improves. Specifically, in almost all tasks, FE provides a better uncertainty quantification with respect to DE, with a lower ECE and NLL. Although FE does not always deliver the highest accuracy, it shows a comparable performance across all tasks. Particularly, it provides the best accuracy in ARC-C under both deterministic (M=0) and stochastic (M=10) settings. It can be concluded from Table E that a more flexible model yields more reliable uncertainty quantification, with no significant loss in predictive accuracy.
To assess the effect of model complexity on generalization, we further compared DE and FE under out-of-distribution scenario, using the models fine-tuned on OBQA from Table E. As Table E indicates, FE consistently delivers superior uncertainty quantification over all tasks under both smaller and larger distribution shifts. It also delivers the best accuracy in almost all tasks.
These findings from Tables E and E confirm that richer stochastic structures enhance both generalization and uncertainty estimation, motivating our focus on contextualizing the full matrix $\mathbf{E}$.
@@
[…]
@@ Appendix G Flipout
12:     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}(\mathbf{x},y)$triangleright Eq. 7
13:     $\varphi\leftarrow\varphi-\eta\nabla_{\varphi}\mathcal{L}(\mathbf{x},y)$triangleright Eq. 8
@@
19:              Save $\boldsymbol{\theta}$, $\varphi$triangleright Checkpoint best model
@@ Appendix H Discussion on Results of Laplace Approximation
For experiments involving fine-tuning via Laplace approximation (LA), we used the publicly released code provided by the authors of [19]. However, the results that we attempted to reproduce—reported in Table 4.2 in our main paper—are substantially worse than those reported in the original paper [19]. This discrepancy is especially pronounced for the OBQA and BoolQ datasets. Our findings are consistent with those reported by the authors of BLoB [20], which suggests that the degradation in performance may stem from sub-optimal MAP solutions that negatively impact the quality of the Laplace approximation-based LoRA fine-tuning.
@@
Appendix I Broader Impacts
@@
Appendix J Additional Information

### v3

@@ 5 Related Works
LLMs have demonstrated remarkable successes across a wide range of applications, including code generation, scientific reasoning, and open-domain question answering; however, they are also notorious for hallucination—a phenomenon in which the model generates fluent but factually incorrect or unsupported content that is not grounded in the input or external knowledge. To mitigate hallucination, a variety of strategies have been proposed, including Reinforcement Learning from Human Feedback (RLHF) to align model outputs with human preferences (liang-etal-2024-learning, ; wen-etal-2025-policy, ), and Retrieval-Augmented Generation (RAG) to ground responses in external knowledge (lewis2020retrieval, ; gao2023retrieval, ). Additionally, probabilistic methods such as BNNs (wang2024blobbayesianlowrankadaptation, ; yang2024bayesianlowrankadaptationlarge, ) and conformal prediction (kumar2023conformal, ; ICLR2024_31421b11, ) have been demonstrated to be able to quantify uncertainty and flag unreliable outputs. Recent work also shows the effectiveness of prompt-level interventions (jiang2023calibrating, ; abbasi-yadkori2024to, ) by carefully crafting instructions and uncertainty-aware prompts to reduce hallucinated content without altering the model architecture.
@@ 6 Conclusion and Discussion
While C-LoRA establishes a promising foundation for uncertainty-aware parameter-efficient fine-tuning, several avenues remain open for future exploration. One interesting direction is extending the framework to multi-modal or multi-task settings, where capturing cross-modal or task-dependent uncertainty could further enhance robustness and generalization. Another promising avenue involves integrating C-LoRA with active learning or adaptive data acquisition strategies, leveraging its uncertainty estimates to guide sample-efficient model updates. Moreover, exploring hierarchical or structured priors within the Bayesian formulation may offer richer representations of both epistemic and aleatoric uncertainty.
@@ 7 Limitations
In this study, we restricted our experiments to fine-tuning LLaMA2-7B models, which limits our understanding of how C-LoRA scales to larger architectures. Investigating the scaling behavior of C-LoRA on larger models remains an open challenge. Furthermore, evaluating uncertainty in language models remains challenging due to the lack of standardized benchmarks with ground-truth uncertainty labels. While our evaluation—based on held-out test sets and established metrics like ECE and NLL—provides meaningful insight into model calibration and predictive confidence, these metrics offer only an indirect view of uncertainty quality. Developing task-specific benchmarks or human-in-the-loop protocols for evaluating uncertainty in NLP would be a valuable direction. Moreover, while C-LoRA is motivated by data-dependent/aleatoric uncertainty, this work does not provide a theoretical guarantee that it can disentangle aleatoric and epistemic uncertainty, which we leave to future investigation.
We also note that, as shown throughout our experiments, C-LoRA overall provides a more calibrated predictions and better uncertainty estimation at the expense of marginally reduced predictive performance which can be preferable in many high-stakes applications. However, in other domains such as spam detection where accuracy is more critical, this trade-off may not be desirable. Hence we acknowledge that this trade-off is task dependent and should be considered before deploying C-LoRA (and in general any uncertainty-aware method).
Finally, our approach employs amortized variational inference to approximate the posterior distribution over the LoRA adapter parameters, providing a scalable and principled Bayesian treatment suitable for large language models. This variational approximation, based on a Gaussian family and a reverse KL objective, is inherently mode-seeking and may underestimate posterior uncertainty or fail to capture multi-modality in the true posterior. Investigating richer variational families and more comprehensive posterior predictive analyses represents an important direction for future work.
8 Acknowledgment
A.H.R., W.Z., Y.W., and X.N.Q. acknowledge the support from United States National Science Foundation (NSF) grants DMREF-2119103, SHF-2215573, and IIS-2212419. S.J., B.J.Y., N.M.U., and X.N.Q. acknowledge the support from the United States Department of Energy’s Office of Science Biological and Environmental Research (BER) program under project B&R# KP1601017 and FWP#CC140. Many of the numerical experiments were conducted using advanced computing resources provided by Texas A&M High Performance Research Computing.
@@ Appendix E On the Importance of Flexibility
Table 9 summarizes the performance of DE and FE with M=0 and M=10 under in-distribution scenario for the six common-sense reasoning tasks. These models are ordered from top to bottom for each metric, such that each successive model is more flexible than the previous one. From Table 9, it is clear that as the flexibility increases, the performance generally improves. Specifically, in almost all tasks, FE provides a better uncertainty quantification with respect to DE, with a lower ECE and NLL. Although FE does not always deliver the highest accuracy, it shows a comparable performance across all tasks. Particularly, it provides the best accuracy in ARC-C under both deterministic (M=0) and stochastic (M=10) settings. It can be concluded from Table 9 that a more flexible model yields more reliable uncertainty quantification, with no significant loss in predictive accuracy.
To assess the effect of model complexity on generalization, we further compared DE and FE under out-of-distribution scenario, using the models fine-tuned on OBQA from Table 9. As Table 10 indicates, FE consistently delivers superior uncertainty quantification over all tasks under both smaller and larger distribution shifts. It also delivers the best accuracy in almost all tasks.
These findings from Tables 9 and 10 confirm that richer stochastic structures enhance both generalization and uncertainty estimation, motivating our focus on contextualizing the full matrix $\mathbf{E}$.
@@
FE (M=0) $\underline{0.89}{\pm_{0.07}}$} $\underline{1.85}\pm_{0.11}$ $\underline{0.85}\pm_{0.05}$} $\underline{0.69}\pm_{0.03}$} $\underline{0.68}\pm_{0.04}$ $\underline{0.34}\pm_{0.00}$
@@ Appendix G Flipout
12:  $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}(\mathbf{x},y)$ $\triangleright$ Eq. 7
13:  $\varphi\leftarrow\varphi-\eta\nabla_{\varphi}\mathcal{L}(\mathbf{x},y)$ $\triangleright$ Eq. 8
@@
19:     Save $\boldsymbol{\theta}$, $\varphi$ $\triangleright$ Checkpoint best model
@@ Appendix H Discussion on Results of Laplace Approximation
For experiments involving fine-tuning via Laplace approximation (LA), we used the publicly released code provided by the authors of [19]. However, the results that we attempted to reproduce—reported in Table 1 in our main paper—are substantially worse than those reported in the original paper [19]. This discrepancy is especially pronounced for the OBQA and BoolQ datasets. Our findings are consistent with those reported by the authors of BLoB [20], which suggests that the degradation in performance may stem from sub-optimal MAP solutions that negatively impact the quality of the Laplace approximation-based LoRA fine-tuning.
@@
Appendix I Discussion on the choice of prior
In this work, we adopt a fixed Gaussian prior which, when paired with a variational Gaussian posterior, enables a closed-form KL divergence and further facilitates performance comparison across baseline methods. However, due to Monte Carlo approximation of the ELBO, alternative richer distributions could also be considered as priors. While richer priors may offer empirical benefits in specific settings, they can introduce additional computational and design challenges. Moreover, due to the highly expressive contextual modules, despite the simplicity, fixed Gaussian prior mainly act as a light regularizer rather than as a strict inductive bottleneck. Prior work [71] has also shown that Bayesian Model Averaging (BMA) is robust to the choice of prior, with posterior predictive behavior remaining similar across different prior families.
Appendix J Empirical Study of Contextualized Variance
To further illustrate that our model captures input-dependent uncertainty, we analyze the predictive variance behavior of the LM head for three OBQA questions, each sharing the same correct label (option “C”). For each question, we extract the predicted variance matrix from the final layer, compute the mean variance across tokens, and report the $\ell_{2}$ norm of the resulting vector to summarize overall uncertainty.
Table 11: Example questions and corresponding variance magnitudes. Despite sharing the same label, the model assigns distinct variances to each question, reflecting input-dependent uncertainty.
Question Variance $\ell_{2}$ norm
Q1: What would light bounce off of when it hits it? 3.7661
Q2: A mother births what? 3.6792
Q3: What happens when animals in hot environments are active? 3.9443
Despite sharing the same label, the model assigns distinct variances to each question, reflecting differences in ambiguity and specificity of the input. This confirms that C-LoRA modulates its predictive distribution based on input semantics, not merely on class labels.
We further examined whether the learned variance meaningfully affects the model’s representations. Specifically, we computed the $\ell_{2}$ norm of the difference between the last-token embedding at $M=0$ and the mean embedding across $M=10$ stochastic samples. For the three representative examples (Q1–Q3), the differences ranged from 6.18 to 7.25, while the embedding norms themselves were approximately 120–122, corresponding to a relative change of about 5–6%. In embedding space, this level of perturbation is non-trivial and indicates meaningful input-dependent variability introduced by the contextualized variance modules.
We extended this analysis to a broader set of examples, comparing 8 correct and 8 incorrect predictions. For each, we computed the $\ell_{2}$ norm of the difference between the deterministic embedding ($M=0$) and the mean of embeddings from $M=10$ samples, then averaged across each group. Correct predictions had an average difference of $6.7\pm 1.09$, while incorrect predictions showed higher variability at $8.45\pm 2.66$. This systematic increase in embedding perturbation for incorrect predictions suggests that C-LoRA’s learned variance corresponds to greater uncertainty for ambiguous or difficult inputs, further supporting its interpretation as modeling input-specific uncertainty.
Appendix K Visualization of Results
For a visual performance comparison, we present the results under in-distribution scenario, after temperature scaling, and the results under out-of-distribution scenario, reported in Tables 1, 2, and 3, using bar plots.
Figure 3 presents bar plots summarizing the results reported in Table 1. As discussed in Section 4.2, all methods exhibit similar predictive performance overall, while C-LoRA and BLoB ($M=10$) achieve superior uncertainty quantification, as measured by ECE and NLL. Figure 3 also illustrates the effect of temperature scaling on calibration, corresponding to Table 2. As noted in Section 4.2, temperature scaling substantially improves calibration. Finally, Figure 4 visualizes the out-of-distribution results from Table 3 via bar plots. Consistent with in-distribution results, performance degrades as the distribution shifts; however, predictive performance remains largely similar across methods, with C-LoRA and BLoB demonstrating more robust uncertainty quantification according to ECE and NLL. As we mentioned earlier, it worth noting that, although C-LoRA offers similar predictive performance overall, it can underperform on certain datasets in exchange for better uncertainty quantification.
Figure 2: Visualization of the results in Table 1.
Figure 3: Visualization of the results in Table 2.
Figure 4: Visualization of the results in Table 3.
Appendix L Broader Impacts
@@ Appendix I Discussion on the choice of prior
Appendix M Additional Information